In the same sample of 300 25-35 year olds, you decide to create a obesity (obese versus not obese). You are interested in examining the relationship between obesity, years of education, and race. 2. dichotomous variable for Write the equation of the multiple regression model with only main effects. (0.5pt) a.
b. Using the following regression table: Obese! Coef. Std. Err. z Pipl p> Izl [95% Conf. Interval) Years Edu I 043 Woman | 141 3.560.000 0.034 -266 1691.58 0. 11a 012 066 068 -019 010 271 598 064 2.12 Intercept . Write out the main effects regression model equation using values for the coefficients. (0.Spt) li. Identify which predictor variable(s) is statistical obesity. Explain why. (1pt) ili. Calculate the odds ratio for each predictor variable. Provide an interpretation of each odds ratio. (1pt) Year edu 0.96 Woman 1.15
Answer #1

a.

Multiple regression models with only main effects is,

ln(P/1-P) = \beta_0 + 61 Years Edu + \beta_2 race

where P is the probability of being obese.

b.

i.

With the given coefficients, the models with only main effects is,

ln(P/1-P) = -0.266 - 0.043 Years Edu + 0.141 Woman

ii.

P-value (P > |z|) for Years Edu is less than 0.05 significance level. Thus, Years Edu is statistically significant with obesity.

P-value (P > |z|) for Women is greater than 0.05 significance level. Thus, Women is not statistically significant with obesity.

iii.

For, Years Edu = 0.96, Women = 1.15

ln(P/1-P) = -0.266 - 0.043 * 0.96 + 0.141 * 1.15 = -0.14513

odds = P/(1-P) = exp(-0.14513) = 0.8649